PHY 2411 University Physics I (Week 10 (Chapters 13 &amp:amp: 14) Assignment Problem 14.4 A 400 m x120 m swimming pool slopes linearly from a 1/20 m depth at one end to a 3.10 m depth at the other Part A What is the mass of water in the pool? Express your answer with the appropriate units
Answer #1

Solution)

Here, volume

V₁ = 4 * 12 * 1.2
V₁ = 57.6 m³

Now, another volume whose depth is 3.10 m

V₂ = .5 * 4 * 12 * 3.1
V₂ = 74.4 m³

Total volume
V = 132 m³

Now, we know

d = m / V

We know, density of water is 1,000 kg per m³

So, m = 1,000(132)

Mass, m = 132,000kg (Ans)

=======

Good luck!:)